OpenStudy (anonymous):

find both solutions to x^2+8x=20

OpenStudy (anonymous):

I think its -10 and 2

OpenStudy (anonymous):

2,-10

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

x^2+8x-20=0 (x+10)(x-2)=0 answer x=-10, or x=+2